OpenStudy (anonymous):

If 2^1.5894=3, then 2^4.5894=?

OpenStudy (tkhunny):

Let's see... 4.5894 / 1.5894 = 2.8875 Well, then I guess \(2^{4.5894} = 3^{2.8875}\) I'm not sure that get's us anywhere. Are you SURE you have the right problem statement and instructions? Maybe, \(2^{4.5894} = 2^{1.5894}\cdot 2^{3}\) That might lead to something.

OpenStudy (anonymous):

That's the right problem

OpenStudy (anonymous):

If it helps the answer choices are 18, 24, 28, 30, 31

OpenStudy (tkhunny):

Okay, now respond to my 2nd theory.

OpenStudy (anonymous):

ya that's correct

OpenStudy (tkhunny):

And you managed what answer?

OpenStudy (anonymous):

The calculator gave me 24.0739

OpenStudy (anonymous):

If we use the value that 2^1.5894 =3 and then multiply 3 x 2^3 the answer is 24

OpenStudy (tkhunny):

Medals 1 We are not concerned with the calculator. \(2^{1.5894}=3.0092417260913488698923493617444\) It's not quite 3. We're going with the premise that \(2^{1.5894}=3\) Now, we have to find \(2^{4.5894}=2^{1.5894+3}=2^{1.5894}⋅2^{3}\). What's your plan?

OpenStudy (tkhunny):

Or, was that it? Since it worked and was PERFECT!!!

OpenStudy (anonymous):

Thanks